OpenStudy (anonymous):

One of the factors of g2 – 32g + 256 is: g + 8 g – 8 g – 16 g + 16

OpenStudy (nurali):

g² - 32g + 256 = 0 g - 16g = - 256 + (- 16)² g - 16g = - 256 + 256 (g - 16)² = 0 Answer: g - 16